# Optimal sampling of tensor networks targeting wave function’s fast decaying tails

### v1

@@ V Conclusions
Suppose we want to sample from the $n$-body quantum state |$\ket{\psi}$ represented as an isometric tensor network Zaletel and Pollmann (2020) with bond dimension $\chi$. In Fig. 9, we report the procedure for an MPS with local dimension $d=2$. In general, the steps (a) to (d) following Fig. 9 are the following:
@@
    where $m,n$ represent the physical indexes of local dimension $d$, while the auxiliary indexes, i.e. indexes not representing the Hilbert space of site 00, are defined as $i_{1},i_{2}\dots i_{n}$. The complex conjugate tensor is indicated by a superscript ***.
@@
The entire procedure scales as $O(n\chi^{2}d^{2})$ for MPS Stoudenmire and White (2010), while as $O(n\chi^{3})$ for TTNs. Since each sample is independent from the others the standard sampling can be trivially parallelized with OPENMP Chandra et al. (2001).
.3.1 Caching the MPS intermediate measurements
Figure 10: Caching for matrix product states (MPS). Example of OPES with $n=3$ sites and local dimension $d=2$. The intervals represent the conditional probabilities of the state represented above it, i.e. the probability $p_{0}(0)=0.5$, $p_{1}(0|p_{0}=0)=0.5$, and so on. First, we compute the probability associated with sample $s_{0}=0.05$ (blue line), obtaining the probability interval $\delta_{0}=(0,0.1)$ associated to the state |$\ket{000}$. We cache the couples $(0,\widetilde{T}_{0})$, $(00,\widetilde{T}_{1})$, where with $\widetilde{T}_{i}$ we indicate the tensor after the projection with the measured basis state. When we continue with the sample $s_{1}=0.29$, we use the cached information and do not perform any operation on $T_{0}$. At the level of $T_{1}$, we go to the right instead of the left; thus, we can eliminate the cached information for $T_{1},T_{2}$. We can do so because any further samples go to the right of the current interval, due to the ascending ordering.
One of the reasons for the speedup of OPES over the standard sampling is the caching of the intermediate states inside a superiteration. We suppose the left-to-right sampling ordering discussed in Sec. .3, and that all $\{T_{j}\}_{0\leq j<m}$, with $m>0$, have been measured. To restart the computation from tensor $T_{m}$, it is sufficient to store the measured state $|b_{0},b_{1},\dots,b_{m-1}\rangle$ and the tensor $T_{m}$ contracted with tensor $T_{m-1}$ after the projection on state $|b_{m-1}\rangle$. The caching in the superiteration is optimal because, due to the ascending ordering of the samples $n_{s}$, we know that once the basis $b_{i}$ of the tensor $T_{i}$ pass from 00 to $1$ we can delete all the cached information for $j<i$ since we completely explored that probability interval. The caching structure is shown in Fig. 10 for an MPS with $n=3$ sites and $d=2$ basis. Performance can be further optimized by keeping the cache between superiterations.
This procedure can be parallelized to using OPENMP Chandra et al. (2001), by assigning to different processors different "branches" of the tree, such as the blue and orange path from tensor $T_{1}$ in Fig. 10. Ensuring an equal workload for each process is more challenging for the OPES than for the standard sampling.
.3.2 Caching the TTN intermediate measurements

### v2

@@ Appendix B Details of the tensor networks implementation
Suppose we want to sample from the $n$-body quantum state |$\ket{\psi}$ represented as an isometric tensor network [61] with bond dimension $\chi$. In Fig. 10, we report the procedure for an MPS with local dimension $d=2$. In general, the steps (a) to (d) following Fig. 10 are the following:
@@
    where $m,n$ represent the physical indexes of local dimension $d$, while the auxiliary indexes, i.e., indexes not representing the Hilbert space of site 00, are defined as $i_{1},i_{2}\dots i_{n}$. The complex conjugate tensor is indicated by a superscript *.
@@
The entire procedure scales as $O(n\chi^{2}d^{2})$ for MPS [83], while as $O(n\chi^{4})$ for TTNs. Since each sample is independent from the others the standard sampling can be trivially parallelized with OPENMP [84].
B.1 Caching the MPS intermediate measurements
Figure 11: Caching for matrix product states (MPS). Example of OPES with $n=3$ sites and local dimension $d=2$. The intervals represent the conditional probabilities of the state represented above it, i.e., the probability $p_{0}(0)=0.5$, $p_{1}(0|p_{0}=0)=0.5$, and so on. First, we compute the probability associated with sample $s_{0}=0.05$ (blue line), obtaining the probability interval $\delta_{0}=(0,0.1)$ associated to the state |$\ket{000}$. We cache the couples $(0,\widetilde{T}_{0})$, $(00,\widetilde{T}_{1})$, where with $\widetilde{T}_{i}$ we indicate the tensor after the projection with the measured basis state. When we continue with the sample $s_{1}=0.29$, we use the cached information and do not perform any operation on $T_{0}$. At the level of $T_{1}$, we go to the right instead of the left; thus, we can eliminate the cached information for $T_{1},T_{2}$. We can do so because any further samples go to the right of the current interval, due to the ascending ordering.
One of the reasons for the speedup of OPES over the standard sampling is the caching of the intermediate states inside a superiteration. We suppose the left-to-right sampling ordering discussed in Sec. B, and that all $\{T_{j}\}_{0\leq j<m}$, with $m>0$, have been measured. To restart the computation from tensor $T_{m}$, it is sufficient to store the measured state $|b_{0},b_{1},\dots,b_{m-1}\rangle$ and the tensor $T_{m}$ contracted with tensor $T_{m-1}$ after the projection on state $|b_{m-1}\rangle$. The caching in the superiteration is optimal because, due to the ascending ordering of the samples $n_{s}$, we know that once the basis $b_{i}$ of the tensor $T_{i}$ pass from 00 to $1$ we can delete all the cached information for $j<i$ since we completely explored that probability interval. The caching structure is shown in Fig. 11 for an MPS with $n=3$ sites and $d=2$ basis. Performance can be further optimized by keeping the cache between superiterations.
This procedure can be parallelized to using OPENMP [84], by assigning to different processors different "branches" of the tree, such as the blue and orange path from tensor $T_{1}$ in Fig. 11. Ensuring an equal workload for each process is more challenging for the OPES than for the standard sampling.
B.2 Caching the TTN intermediate measurements